OpenStudy (anonymous):

You are buying $20 worth of bird seed that consists of two types of seed. Thistle seed costs $2 a pound and dark oil seed costs $1.50 a pound. Write an equation in standard form that represents the different amounts of $2 thistle seed, x, and $1.50 dark oil seed, y, that you could buy.

OpenStudy (anonymous):

I know that I would clear the decimals to get 200x + 150y = 2000. now is that my answer?

OpenStudy (anonymous):

If you were to write it with decimals, you don't have the chance of making any mistakes.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Instead of multiplying it out by 100, just leave it in decimal format.

OpenStudy (anonymous):

but the numbers are suppose to be integers

OpenStudy (anonymous):

Doesn't specify such in the question, but I get your point.

OpenStudy (anonymous):

It would look like 2x+1.5y=20. Just multiply it by 2 so, you'll get 4x+3y=40

OpenStudy (anonymous):

I just don't see the point of multiplying by 100. Does that help? And yes, the equation that you had would be the correct answer.

OpenStudy (anonymous):

but if i do multiply by 100 and get 200x + 300y = 2000, is it true that the only factor they should share is 1?

OpenStudy (anonymous):

That is not true.

OpenStudy (anonymous):

200x + 150y = 2000. Multiply everything by 2 400x+300y=4000 Now, divide everything by 100.

OpenStudy (anonymous):

why muliply by 2?

OpenStudy (anonymous):

that's where I'm getting lost

OpenStudy (anonymous):

You can multiply by anything.

OpenStudy (anonymous):

By 10, 15, 16 etc. All what you are doing is scaling it by that number.

OpenStudy (anonymous):

thanks